OpenStudy (anonymous):

what are the solutions of Ix+6I greater than or equal to 5? write solutions as either the union or the intersection of two sets

hartnn (hartnn):

if |y|>=a then \[-a \ge y \ge a\]

hartnn (hartnn):

here y=x+6, a=5

OpenStudy (anonymous):

thank you! your great at this

hartnn (hartnn):

welcome ^_^ if you practice, you'll also be great :)